### v1

@@ 4.9 Encoding model of neural representations in the brain
We used the fMRI dataset from the THINGS-data collection [43] to explore whether LLM conceptual representations can map onto brain activity patterns associated with visually grounded concepts. The dataset encompasses brain imaging data from three participants exposed to 8,740 representative images of 720 concepts over 12 sessions. The concepts are sampled from the total of 1,854 concepts in THINGS and are thus relevant to the behavioural data that can be predicted by LLM conceptual representations. We obtained the neural representation of each concept for each participant by averaging over its corresponding images.
To investigate the relationship between LLM-derived conceptual representations and neural representations, we trained a linear encoding model to predict voxel activations based on the LLM-derived representations. Specifically, the activation at each voxel was modelled by $y_{v,c}=\mathbf{h}_{c}\mathbf{w}+b+\epsilon$, where $y_{v,c}$ is the activation at voxel $v$ for concept $c$, $\mathbf{h}_{c}$ is the corresponding LLM representation, $\mathbf{w}$ is a vector of regression coefficients, $b$ is a constant, and $\epsilon$ represents residual error. The parameters $\mathbf{w}$ and $b$ were estimated via regularized linear regression, minimizing the least-squares error with a ridge penalty controlled by hyperparameter $\lambda$, selected through cross-validation within the training set. Model performance was assessed through twenty-fold cross-validation with non-overlapping concepts, evaluating the correlation between predicted and observed neural responses across folds. To establish statistical significance, we generated a null distribution of randomized correlation values from 10,000 test data permutations [107]. Voxel-wise $P$-values were computed by comparing predicted correlations to the null distribution and adjusted for multiple comparisons with FDR correction. We also computed noise-ceiling-normalized $R^{2}$ for each voxel by dividing the original $R^{2}$ by the estimated noise ceiling. The noise ceilings were derived based on signal and noise variance, estimated from neural activity variability to presentations of the same concept [108] (Fig. S12).
@@ 4.10 Variance partitioning between conceptual and alternative representations
We combined each baseline representation with LLM-derived conceptual representations and analysed the shared and unique variance each could account for [107]. To isolate unique variance, we orthogonalized the target representation and the neural responses with respect to the alternate representation, thereby removing the shared variance from both the representation and the fMRI data. The residuals of the target representation were then used to predict the fMRI residuals, and the unique variance explained by the target representation was calculated as the $R^{2}$ using twenty-fold cross-validation. For shared variance, we first concatenated both representations, used them to predict neural responses, and calculated the $R^{2}$ in the same twenty-fold cross-validation to determine the total variance explained. Shared variance was subsequently estimated by subtracting the unique contributions of each representation from this total. For our results, we focused on voxels with a noise ceiling greater than $5\%$ and reported the noise-ceiling-normalized $R^{2}$.

### v3

@@ 4.9 Encoding model of neural representations in the brain
We used the fMRI dataset from the THINGS-data collection [40] to explore whether LLM conceptual representations can map onto brain activity patterns associated with visually grounded concepts. The dataset encompasses brain imaging data from three participants exposed to 8,740 representative images of 720 concepts over 12 sessions. The concepts are sampled from the total of 1,854 concepts in THINGS and are thus relevant to the behavioral data that can be predicted by LLM conceptual representations. We obtained the neural representation of each concept for each participant by averaging over its corresponding images.
To investigate the relationship between LLM-derived conceptual representations and neural representations, we trained a linear encoding model to predict voxel activations based on the LLM-derived representations. Specifically, the activation at each voxel was modeled by $y_{v,c}=\mathbf{h}_{c}\mathbf{w}+b+\epsilon$, where $y_{v,c}$ is the activation at voxel $v$ for concept $c$, $\mathbf{h}_{c}$ is the corresponding LLM representation, $\mathbf{w}$ is a vector of regression coefficients, $b$ is a constant, and $\epsilon$ represents residual error. The parameters $\mathbf{w}$ and $b$ were estimated via regularized linear regression, minimizing the least-squares error with a ridge penalty controlled by hyperparameter $\lambda$, selected through cross-validation within the training set. Model performance was assessed through twenty-fold cross-validation with non-overlapping concepts, evaluating the correlation between predicted and observed neural responses across folds. To establish statistical significance, we generated a null distribution of randomized correlation values from 10,000 test data permutations [104]. Voxel-wise $P$-values were computed by comparing predicted correlations to the null distribution and adjusted for multiple comparisons with FDR correction. We also computed noise-ceiling-normalized $R^{2}$ for each voxel by dividing the original $R^{2}$ by the estimated noise ceiling. The noise ceilings were derived based on signal and noise variance, estimated from neural activity variability to presentations of the same concept [105] (Fig. S19).
@@ 4.10 Variance partitioning between conceptual and alternative representations
We combined each baseline representation with LLM-derived conceptual representations and analyzed the shared and unique variance each could account for [104]. To isolate unique variance, we orthogonalized the target representation and the neural responses with respect to the alternate representation, thereby removing the shared variance from both the representation and the fMRI data. The residuals of the target representation were then used to predict the fMRI residuals, and the unique variance explained by the target representation was calculated as the $R^{2}$ using twenty-fold cross-validation. For shared variance, we first concatenated both representations, used them to predict neural responses, and calculated the $R^{2}$ in the same twenty-fold cross-validation to determine the total variance explained. Shared variance was subsequently estimated by subtracting the unique contributions of each representation from this total. For our results, we focused on voxels with a noise ceiling greater than $5\%$ and reported the noise-ceiling-normalized $R^{2}$.
[…]
Supplementary information
The supplementary information contains supplementary text, Figures S8 to S19, and Tables S1 to S2.
[…]
Data and Code Availability
This work makes use of previously published datasets [42, 40, 95, 47, 49, 46]. The code and additional data for reproducing all our experiments will be made publicly accessible upon publication on GitHub and Zenodo.